Company A has a beta of 1.5. Company B has a beta of 0.90. The required rate of return on the stock market is 10.5%. The risk-free rate is 5%. How much greater is Company A's required rate of return compared to Company B's?
Answer #1

required return= risk-free rate +Beta*(market rate- risk-free rate )

required return for A=5+1.5*(10.5-5)=13.25%

required return for B=5+0.9*(10.5-5)=9.95%

Hence excess required return=(13.25-9.95)%

=3.3%